OpenStudy (anonymous):

Solve 2x-7y=13 for y. This is on a final and I do not know where to begin.

OpenStudy (anonymous):

2x-7y=13 subtract 2x from both sides -7y=13-2x divide both sides by -7 y=(-13/7)+(2x)/7

OpenStudy (anonymous):

To get y by itself look on the left side of the = sign and you see 2x - 7y you need to get the y by itself, so the first thing you get rid of is 2x. To get rid of 2x you subtract 2x so 2x - 2x - 7y = 13 - 2x adding like terms -7y = 13 - 2x look at the 7 term again and we need to get rid of the -7 the -7 is multiplied by y so we need to divide by -7 -7y = 13 - 2x --- ------ -7 -7 so y = 13 - 2x ------- -7

OpenStudy (anonymous):

got it